A cord is wrapped around a spool with radius 20 cm
Answer #1

here,

radius of the spool , r = 0.2 m

v = v0 + k0 * t^2

v0 = 0.05 m/s

k0 = 0.001 m/s^3

the angular velocity , w = v/r

w = (v0 + k0 * t^2)/r

and

differentiating the equation of angular accelration

angular accelration , alpha = ( 0 + 2 * k0 * t ) /r

alpha = 2 * k0 * t /r